# OmniSAM: Omnidirectional Segment Anything Model for UDA in Panoramic Semantic Segmentation

### v1

@@ Abstract
Segment Anything Model 2 (SAM2) has emerged as a strong base model in various pinhole imaging segmentation tasks. However, when applying it to $360^{\circ}$ domain, the significant field-of-view (FoV) gap between pinhole ($70^{\circ}\times 70^{\circ}$) and panoramic images ($180^{\circ}\times 360^{\circ}$) poses unique challenges. Two major concerns for this application includes 1) inevitable distortion and object deformation brought by the large FoV disparity between domains; 2) the lack of pixel-level semantic understanding that the original SAM2 cannot provide. To address these issues, we propose a novel OmniSAM framework, which makes the first attempt to apply SAM2 for panoramic semantic segmentation. Specifically, to bridge the first gap, OmniSAM first divides the panorama into sequences of patches. These patches are then treated as image sequences in similar manners as in video segmentation tasks. We then leverage the SAM2’s memory mechanism to extract cross-patch correspondences that embeds the cross-FoV dependencies, improving feature continuity and the prediction consistency along mask boundaries. For the second gap, OmniSAM fine-tunes the pretrained image encoder and reutilize the mask decoder for semantic prediction. An FoV-based prototypical adaptation module with dynamic pseudo label update mechanism is also introduced to facilitate the alignment of memory and backbone features, thereby improving model generalization ability across different sizes of source models. Extensive experimental results demonstrate that OmniSAM outperforms the state-of-the-art methods by large margins, e.g., 79.06% (10.22$\uparrow$) on SPin8-to-SPan8, 62.46% (6.58$\uparrow$) on CS13-to-DP13.
Figure 1: Performance comparison between our OmniSAM and SoTA methods [39, 42].
@@ 1 Introduction
The expansive field of view (FoV) of $360^{\circ}\times 180^{\circ}$ brought by the panoramic cameras has led to their growing popularity in diverse applications, such as autonomous driving and virtual reality [7, 1, 9]. Unlike pinhole cameras, which are limited by a confined FoV, panoramic imaging enables a more holistic perception of the environment. Consequently, recent research efforts focus on leveraging this advantage to enhance scene understanding for intelligent systems. Generally, panoramic cameras capture $360^{\circ}$ visual data in a spherical format, which is projected onto a 2D planar representation to facilitate compatibility with conventional imaging pipelines. Unlike pinhole images, this projection process introduces inevitable distortions and object deformations due to the non-uniform distribution of pixels. Moreover, developing effective panoramic segmentation models is challenging due to the scarcity of large-scale, precisely annotated datasets, as panoramic image annotation remains a labor-intensive task. For these reasons, unsupervised Domain Adaptation (UDA) for panoramic semantic segmentation [36, 42, 39, 40, 41, 13] has been proposed to bridge the domain gap between pinhole and the panoramic image domain.
@@ 2.2 UDA for Panoramic Semantic Segmentation
Domain adaptation enhances a model’s generalization ability to unseen domains and has been widely applied in panoramic semantic segmentation [40, 39, 42, 41, 35, 36, 37, 38]. It can be categorized into three types: adversarial learning [11, 27, 3], pseudo-labeling [37, 38], and prototypical adaptation [40, 42, 41, 39, 35, 36, 34]. Adversarial training captures domain-invariant characteristics by using a discriminator to force the target model to generate indistinguishable features across domains. Pseudo-labeling generates self-supervised labels for the target domain. GoodSAM [37] and GoodSAM++ [38] refine these labels using SAM, providing additional knowledge to the target model. The prototypical approach aligns high-level feature centers between domains. Previous methods [35, 37, 36, 38, 40, 39, 42, 41] focus on semantic details, and distortion in panoramic FoV during prototype extraction. However, these methodologies treat the equirectangular projection (ERP) images or tangent projection (TP) as pinhole images and forward the entire images to the model with no specific designs, causing significant prototypical gap across domains. To better accommodate the panoramas for SAM2 and efficient domain adaptation, we takes a totally different perspective on processing the target panoramas by generating overlapping FoV sequences of patches with a sliding window cropping strategy. Consequently, a FPA module is then introduced for cross-domain prototypical adaptation with patches of various FoV.
@@ 3.4 FoV-based Prototypical Adaptation (FPA)
As preliminaries for domain adaptation, we first define the segmentation loss computed individually for the source and target domains. Given $\{x^{S},y^{S}\}$ and $x^{T}$, we forward them into the pretrained source model $F^{S}$:
@@
OmniSAM-S 77.13 56.12 (+4.00) -21.52
OmniSAM-B 77.00 56.49 (+5.09) -20.51
OmniSAM-L 79.50 56.61 (+5.21) -22.89
@@ 4.1 Experimental Setup
Preprocessing: All models are trained on 2 NVIDIA GPUs with an initial learning rate of $6\times 10^{-5}$, scheduled using a polynomial decay strategy with a power of $0.9$. We use the AdamW optimizer with epsilon of $10^{-8}$ and weight decay of $10^{-4}$. The batch size is set to 4 for OmniSAM-L and OmniSAM-B, and 6 for OmniSAM-S and OmniSAM-T per GPU. The memory bank size is set to 9. We use a sliding window to crop the rectangular image, resulting in a 9-frame FoV sequence of $1024\times 1024$ images. The sliding stride depends on the resolution of the images. The SPin dataset provides $1080\times 1080$ images, which are resized to $1024\times 1024$ and do not require sliding. Meanwhile, the stride is set to $128$ for both the CS and SP datasets with $2048\times 1024$ images. The SPan dataset contains $4096\times 2048$ images, which include inherent black regions brought by equirectangular projection. Hence we remove these redundant areas by cropping the images and downscale them to $3072\times 1024$. The corresponding sliding stride is $256$. The DensePASS dataset contains $2048\times 400$ images, which are resized to $4096\times 1024$, with a sliding stride of $384$.
Trainable Parameters: We evaluate the trainable parameters of our OmniSAM model by examining the impact of adding or removing the memory attention module. The parameter sizes for different variants of the OmniSAM model are presented in the suppl. mat.. Without memory attention, fine-tuning via LoRA-based adaptation affects only a small subset of the model, resulting in minimal trainable parameters across all variants. The smallest (OmniSAM-T) contains 1.45 MB, while the largest (OmniSAM-L) has 2.81 MB. In contrast, the memory attention module contributes approximate 22.60 MB parameters. Specifically, OmniSAM-T increases from 1.45 MB to 24.05 MB, and OmniSAM-L grows from 2.81 MB to 25.41 MB.
@@ 4.3 Experimental Results
Next, we evaluate our framework in the CS13-to-DP13 setting. As shown in Table 3, with adaptation incorporated, all variants surpass the DTA4PASS model [13], which is based on multiple source datasets, achieving improvement in mIoU of $3.73\%$ to $6.58\%$. Even without memory modules, OmniSAM-S, OmniSAM-B and OmniSAM-L outperform the baseline DATR-S, which employs CFA adaptation [39]. The visualization results in Fig. 5 also demonstrate the superiorty of our OmniSAM. Additionally, we explore the capability of our model within the Syn2Real scenario. The quantitative results of SP13-to-DP13 is presented in Table 4. The OmniSAM-B yields $0.56\%$ mIoU improvement. More quantitative comparison and visualization results refer to the suppl. mat..
@@ 5 Ablation Study
Ablation of $\lambda$. Finally we conduct experiments on OmniSAM-B in the outdoor scenario to evaluate the impact of $\lambda$. As shown in Table 6, OmniSAM-B achieves the highest score in mIoU while $\lambda=0.1$.
$\lambda$ 0 0.01 0.1 0.2 0.5 1.0
mIoU 56.49 61.64 62.46 61.59 61.44 60.90
$\Delta$ +5.15 +5.97 +5.10 +4.95 +4.41
Table 6: Ablation Study of $\lambda$.
@@ 1 Supplementary Implementation Details
Training and Adaptation Details. During training, We introduce randomness to the window’s sliding direction, including forward sliding (left-to-right) and reverse sliding (right-to-left). This helps model learn more cross-patch dependencies. For target domains in our domain adaptation, the SPan dataset contains $4096\times 2048$ images, which include inherent black regions brought by equirectangular projection. Hence we remove these redundant areas by cropping the images and downscale them to $3072\times 1024$. The corresponding sliding stride is $256$. The DensePASS dataset contains $2048\times 400$ images, which are resized to $4096\times 1024$, with a corresponding sliding stride of $384$. Besides, we sample 400 images from the target domain for pseudo-labels updating in each epoch.
@@ 2.1 Trainable Parameters
Table 7 presents the detailed trainable parameters of our proposed OmniSAM framework. The LoRA fine-tuning on the pre-trained image encoder only requires less than 3MB parameters. The full fine-tuning on memory attention module requires training on approximate 22.6 MB parameters.
@@
     OmniSAM-T w/o MA      1.45
     OmniSAM-S w/o MA      1.54
     OmniSAM-B w/o MA      1.81
     OmniSAM-L w/o MA      2.81
     OmniSAM-T w/ MA      24.05
     OmniSAM-S w/ MA      24.14
     OmniSAM-B w/ MA      24.41
     OmniSAM-L w/ MA      25.41

### v2

@@ Abstract
Segment Anything Model 2 (SAM2) has emerged as a strong base model in various pinhole imaging segmentation tasks. However, when applying it to $360^{\circ}$ domain, the significant field-of-view (FoV) gap between pinhole ($70^{\circ}\times 70^{\circ}$) and panoramic images ($180^{\circ}\times 360^{\circ}$) poses unique challenges. Two major concerns for this application includes 1) inevitable distortion and object deformation brought by the large FoV disparity between domains; 2) the lack of pixel-level semantic understanding that the original SAM2 cannot provide. To address these issues, we propose a novel OmniSAM framework, which makes the first attempt to apply SAM2 for panoramic semantic segmentation. Specifically, to bridge the first gap, OmniSAM first divides the panorama into sequences of patches. These patches are then treated as image sequences in similar manners as in video segmentation tasks. We then leverage the SAM2’s memory mechanism to extract cross-patch correspondences that embeds the cross-FoV dependencies, improving feature continuity and the prediction consistency along mask boundaries. For the second gap, OmniSAM fine-tunes the pretrained image encoder and reutilize the mask decoder for semantic prediction. An FoV-based prototypical adaptation module with dynamic pseudo label update mechanism is also introduced to facilitate the alignment of memory and backbone features, thereby improving model generalization ability across different sizes of source models. Extensive experimental results demonstrate that our method outperforms the state-of-the-art methods by large margins, e.g., 79.06% (10.22$\uparrow$) on SPin8-to-SPan8, 62.46% (6.58$\uparrow$) on CS13-to-DP13.
@@ 1 Introduction
The expansive field of view (FoV) of $360^{\circ}\times 180^{\circ}$ brought by the panoramic cameras has led to their growing popularity in diverse applications, such as autonomous driving and virtual reality [7, 1, 9, 41]. Unlike pinhole cameras, which are limited by a confined FoV, panoramic imaging enables a more holistic perception of the environment. Consequently, recent research efforts focus on leveraging this advantage to enhance scene understanding for intelligent systems. Generally, panoramic cameras capture $360^{\circ}$ visual data in a spherical format, which is projected onto a 2D planar representation to facilitate compatibility with conventional imaging pipelines. Unlike pinhole images, this projection process introduces inevitable distortions and object deformations due to the non-uniform distribution of pixels. Moreover, developing effective panoramic segmentation models is challenging due to the scarcity of large-scale, precisely annotated datasets, as panoramic image annotation remains a labor-intensive task. For these reasons, unsupervised Domain Adaptation (UDA) for panoramic semantic segmentation [38, 45, 42, 43, 44, 14] has been proposed to bridge the domain gap between pinhole and the panoramic image domain.
@@ 2.2 UDA for Panoramic Semantic Segmentation
Domain adaptation enhances a model’s generalization ability to unseen domains and has been widely applied in panoramic semantic segmentation [43, 42, 45, 44, 37, 38, 39, 40]. It can be categorized into three types: adversarial learning [11, 29, 3], pseudo-labeling [39, 40], and prototypical adaptation [43, 45, 44, 42, 37, 38, 36]. Adversarial training captures domain-invariant characteristics by using a discriminator to force the target model to generate indistinguishable features across domains. Pseudo-labeling generates self-supervised labels for the target domain. GoodSAM [39] and GoodSAM++ [40] refine these labels using SAM, providing additional knowledge to the target model. The prototypical approach aligns high-level feature centers between domains. Previous methods [37, 39, 38, 40, 43, 42, 45, 44, 18, 13] focus on semantic details, and distortion in panoramic FoV during prototype extraction. However, these methodologies treat the equirectangular projection (ERP) images or tangent projection (TP) as pinhole images and forward the entire images to the model with no specific designs, causing significant prototypical gap across domains. To better accommodate the panoramas for SAM2 and efficient domain adaptation, we takes a totally different perspective on processing the target panoramas by generating overlapping FoV sequences of patches with a sliding window cropping strategy.
@@ 3.4 FoV-based Prototypical Adaptation (FPA)
As preliminaries for domain adaptation, we first define the segmentation loss computed individually for the source and target domains. Given $\{x^{S},y^{S}\}$ and $x^{T}$, we forward them into the pretrained source model and target model $F^{T}$ initialized with pretrained source weights:
@@
OmniSAM-S 77.13 56.12 (+4.64) -21.52
OmniSAM-B 77.00 56.49 (+5.01) -20.51
OmniSAM-L 79.50 56.61 (+5.13) -22.89
@@ 4.1 Experimental Setup
Preprocessing: All models are trained on 2 NVIDIA A800 GPU with an initial learning rate of $6\times 10^{-5}$, scheduled using a polynomial decay strategy with a power of $0.9$. We use the AdamW optimizer with epsilon of $10^{-8}$ and weight decay of $10^{-4}$. The batch size is set to 4 for OmniSAM-L and OmniSAM-B, and 6 for OmniSAM-S and OmniSAM-T per GPU. The memory bank size is set to 9. We use a sliding window to crop the rectangular image, resulting in a 9-frame FoV sequence of $1024\times 1024$ images. The sliding stride depends on the resolution of the images. The SPin dataset provides $1080\times 1080$ images, which are resized to $1024\times 1024$ and do not require sliding. Meanwhile, the stride is set to $128$ for both the CS and SP datasets with $2048\times 1024$ images. The SPan dataset contains $4096\times 2048$ images, which include inherent black regions brought by equirectangular projection. Hence we remove these redundant areas by cropping the images and downscale them to $3072\times 1024$. The corresponding sliding stride is $256$. The DensePASS dataset contains $2048\times 400$ images, which are resized to $4096\times 1024$, with a sliding stride of $384$.
Model Parameters: The total parameters and computation costs of our OmniSAM variants are presented in the suppl. mat.. The Tiny and Small variants exhibit competitive parameter counts, similar to state-of-the-art methods. While the Large (209.2M) and the Base (72.0M) variants significantly increase the model size, which may limit its real-time application. A trade-off in inference speed should be considered while choosing the model for specific task.
@@ 4.3 Experimental Results
Next, we evaluate our framework in the CS13-to-DP13 setting. As shown in Table 3, with adaptation incorporated, all variants surpass the DTA4PASS model [14], which is based on multiple source datasets, achieving improvement in mIoU of $3.13\%$ to $6.58\%$. Even without memory modules, OmniSAM-S, OmniSAM-B and OmniSAM-L outperform the baseline DATR-S, which employs CFA adaptation [42]. The visualization results in Fig. 4 also demonstrate the superiorty of our OmniSAM. Additionally, we explore the capability of our model within the Syn2Real scenario. The quantitative results of SP13-to-DP13 is presented in Table 4. The OmniSAM-B yields $0.56\%$ mIoU improvement. More quantitative comparison and visualization results refer to the suppl. mat..
@@ 6 Conclusion
7 Acknowledgement
This work was supported by the Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); CAAI-Ant Group Research Fund; Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
@@ 1 Supplementary Implementation Details
Training and Adaptation Details. During training, We introduce randomness to the window’s sliding direction, including forward sliding (left-to-right) and reverse sliding (right-to-left). This helps model learn more cross-patch dependencies. For target domains in our domain adaptation, the SPan dataset contains $4096\times 2048$ images, which include inherent black regions brought by equirectangular projection. Hence we remove these redundant areas by cropping the images and downscale them to $3072\times 1024$. The corresponding sliding stride is $256$. The DensePASS dataset contains $2048\times 400$ images, which are resized to $3072\times 1024$, with a corresponding sliding stride of $256$. Besides, we sample 400 images from the target domain for pseudo-labels updating in each epoch.
@@ 2.1 Trainable Parameters
We evaluate the trainable parameters of our OmniSAM model by examining the impact of adding or removing the memory attention module. The parameter sizes for different variants of the OmniSAM model are presented in the Table 6. Without memory attention, fine-tuning via LoRA-based adaptation affects only a small subset of the model, resulting in minimal trainable parameters across all variants. The smallest (OmniSAM-T) contains 0.36 MB, while the largest (OmniSAM-L) has 0.70 MB. In contrast, the memory attention module contributes approximate 5.65 MB parameters. Specifically, OmniSAM-T increases from 0.36 MB to 6.01 MB, and OmniSAM-L grows from 0.70 MB to 6.35 MB.
@@
     OmniSAM-T w/o MA      0.36
     OmniSAM-S w/o MA      0.39
     OmniSAM-B w/o MA      0.45
     OmniSAM-L w/o MA      0.70
     OmniSAM-T w/ MA      6.01
     OmniSAM-S w/ MA      6.04
     OmniSAM-B w/ MA      6.10
     OmniSAM-L w/ MA      6.35
@@
2.2 Total Parameters and Computation Costs
The total parameters and computation costs of our OmniSAM variants are presented in the Table 7. The Tiny and Small variants exhibit competitive parameter counts, similar to state-of-the-art methods. While the Large (209.2M) and the Base (72.0M) variants significantly increase the model size, which may limit its real-time application. A trade-off in inference speed should be considered while choosing the model for specific task.
     Network      Param. (M)      FLOPs (G)
     Trans4PASS+-S      44.9      251.1
     DATR-S      25.8      139.2
     360SFUDA++      28.7      148.0
     OmniSAM-T w/ MA      32.0      118.8
     OmniSAM-S w/ MA      38.8      149.8
     OmniSAM-B w/ MA      72.0      280.8
     OmniSAM-L w/ MA      209.2      828.0
Table 7: Computatioal costs for networks with $1024^{2}$ resolution input.
2.3 Ablation of $\lambda$.
We conduct experiments on OmniSAM-B in the outdoor scenario to evaluate the impact of $\lambda$. As shown in Table 8, OmniSAM-B achieves the highest score in mIoU while $\lambda=0.1$.
$\lambda$ 0 0.01 0.1 0.2 0.5 1.0
mIoU 56.49 61.64 62.46 61.59 61.44 60.90
$\Delta$ +5.15 +5.97 +5.10 +4.95 +4.41
Table 8: Ablation Study of $\lambda$.